NASA is expected to send a 2600-kg satellite 450 km above the earth’s surface. (a) What is its radius? (b) What speed will it have? (Hint: Earth’s mass is 5.97 x 1024 kg)


1
Expert's answer
2021-11-29T11:37:19-0500

(a) Radius of the orbit:


"a=R_E+H=6375+450=6825\text{ km}."

(b) Speed (from equilibrium of forces):


"v=\sqrt{GM/a}=7.640\text{ km/s}."